1. Multiply these fractions and reduce to lowest terms: 250/500 X 1/2. Express answer with a number with a decimal point.

2. Digoxin (Lanoxin) toxicity may be characterized by bradycardia, loss of appetite, and yellow halos appearing around objects. TRUE or FALSE?

3. Before beginning therapy with lipid-lowering drugs, the patient must have baseline testing done for ________________________ function.

a. Kidney

b. Cardiac

c. Brain

d. Liver
Answer #1

1. 250/500 can reduce to 1/2

So, we have to multiply 250/500* 1/2

ie, 1/2× 1/2 = (1×1)/(2×2)

= 1/4 = 0.25

2. Digoxin toxicity shows cardiac , gastrointestinal, neurological and visual symptoms.

Cardiac- palpitations, edema on lower extremities, shortness of breath, hypotension, bradycardia, dyspnea

Visual- yellow hallows, photophobia and photopsia, snowy vision

CNS- drowsiness,fatigue, confusion, hallucinations, paresthesia etc.

GI- nausea, vomiting, decreased appetite, weight loss, diarrhoea , abdominal pain

Answer: true

3.lipid lowering drugs are medicine used to reduce low density cholesterol and triglycerides , which increases cardiac risks in its increased state.

It is mainly given to those having high cholesterol would not be controlled only with food or lifestyle changes.so before commencing any lipid lowering drugs, should asses the baseline cardiac function.

Answer: b